A valószínűségszámítás II. gyakorlaton való részvétel

dátum:  IX. 5. IX. 12. IX. 19. IX. 26. X. 3. X.10. X. 17.

diákok:

Bata Lászlő  +  +  +  +  +  +  +
Bordi Gergely  +  −  +  +  +  −
Éliás Anikó  +  +  +  −  +  −  −
Fehér Marianna  +  +  +  +  +  +  −
Funk Evelin  −  −  +  +  +  −  +
Glausits Tamás  +  +  −  +  +
Gyenge Timea  +  −  −  +  +  −
Gyöngy Imre  +  +  −  +  +  +
Herka Erzsébet  +  +  +  −
Juhász Tibor  +  +  +  −  +  +  +
Homoki István  +  −  +  −  +  +
Kabai Zsuzsanna  +  +  +  +  −  −  −
Kaposi Georgina  +  +  −  +  +  +  −
Kávrán Péter  +  −  +  +  +
Keserű Zsuzsanna +  −  +  +  +  +  +
Kozma Ágnes  +  +  +  +  +  +  +
László Anikó  +  −  −  −  +  +  +
Majoros Livia  +  −  +  +  +  +  +
Milo N. Péter  +  +  +  +  +  +
Nagylaki Melinda  +  +  +  −  +
Novák Tamás  +  +  −  +  −
Pável Erika  +  +  +  +  +  −
Ragány Szilvia  +  +  +  +  +  −
Sáfrány Erika  +  −  +  +  +  +  +
Sali Krisztina  +  +  +  −  +  −
Sereg Edgár  +  +  −  +  +  −  +
Sulina Tamás  +  +  +  −  +  −
Szabó Erika  +  +  +  +  −  −
Varga Bernadett  +  +  +  +  +  +  +
Varga Zsuzsanna  +  +  +  +  −
Wittner Ferenc  +  +  +  −  +  +

dátum:  X. 31. XI. 7. XI. 14. XI. 21. XI. 28. XII. 5. XII. 12. XII. 19.

diákok:

Bata Lászlő  +  +  +  +
Bordi Gergely  +  +  +  +
Éliás Anikó  +  +  +  +
Fehér Marianna  −  +  +  +
Funk Evelin  +  −  +  +
Glausits Tamás  +  +  +  −
Gyenge Timea  +  +  +  +
Gyöngy Imre  +  −  +  +
Herka Erzsébet  −  −  +  +
Juhász Tibor  +  −  +  +
Homoki István  +  +  +  +
Kabai Zsuzsanna  +  −  −  +
Koczó Kinga  +  +  +  +
Kaposi Georgina  +  +  −  +
Kávrán Péter  −  +  +  +
Keserű Zsuzsanna  −  +  +  +
Kozma Ágnes  +  +  −  +
László Anikó  +  −  −  +
Majoros Livia  +  +  +  +
Milo N. Péter  +  +  +  +
Nagylaki Melinda  +  +  +  −
Novák Tamás  +  +  +  +
Pável Erika  +  +  +  −
Ragány Szilvia  +  +  +  +
Sáfrány Erika  +  +  +  +
Sali Krisztina  +  +  −  +
Sereg Edgár  −  −  −  +
Sulina Tamás  +  +  −  +
Szabó Erika  +  +  +  +
Varga Bernadett  +  +  −  +
Varga Zsuzsanna  +  +  +  +
Wittner Ferenc  +  +  −  +